Ginny’s Restaurant Problem

Ginny is endowed with $8 million and is deciding whether to invest in a restaurant. Assume perfect capital markets with an interest rate of 6%.

Investment Option

Investment (millions)

End of Year 1 CFs (millions)

End of Year 2 CFs (millions)

1

2

1.8

1.8

2

3

4.3

1.0

3

4

5.4

1.4

4

5

5.2

1.6

  1. List 4 perfect capital market assumptions.
  2. Which investment option should Ginny choose?
  3. Which investment option can be eliminated from consideration? Why?

Ginny is actively pursuing another business venture as a ticket scalper. She estimates that for a $2 million investment in inventory she can resell her tickets for $6 million over the next two years (cash flows realized in exactly two years). Assume the same 6% interest rate.

  1. What is the NPV of the Ticket Brokering venture?
  2. What is the new value of Ginny’s Corporation?
  3. Suppose Ginny does not have the $2 million to start the new venture. Instead, she wants to raise equity capital by issuing 100,000 shares. What price will new investors be willing to pay?
Answer #1

The Solution to Question 1 part i)

The Perfect Capital Market Assumptions are as follows:

  1. There is No transaction cost or flotation cost
  2. Investors are perfectly rational
  3. The Securities are infinitely divisible
  4. Tax-free world

The Solution to Question 1 part ii)

In order to decide which investment option to choose, we will calculate the Net Present Value of all the options:

Investment Option 1:

Year Cash Flows in millions PV factor @ 6% Total PV in millions Cumulative Cash flows
0 (2.00) 1.0 (2.00) (2.00)
1 1.80 0.9434 1.70 (0.20)
2 1.80 0.8900 1.60 1.60
NPV $ 1.30

Pay Back Period = 2 + 0.20/1.80 = 2.11 years

Profitability Index = PV of Cash Inflows / PV of Cash Outflows = 3.30 / 2 = 1.65

Investment Option 2:

Year Cash Flows in millions PV factor @ 6% Total PV in millions Cumulative Cash flows
0 (3.00) 1.0 (3.00) (3.00)
1 4.30 0.9434 4.06 1.30
2 1.80 0.8900 0.89 3.10
NPV $ 1.95

Pay Back Period = 1 + 3/4.30 = 1.70 years

Profitability Index = PV of Cash Inflows / PV of Cash Outflows = 4.95 / 3 = 1.65

Investment Option 3:

Year Cash Flows in millions PV factor @ 6% Total PV in millions Cumulative Cash flows
0 (4.00) 1.0 (4.00) (4.00)
1 5.40 0.9434 5.09 1.40
2 1.40 0.8900 1.25 2.80
NPV $ 2.34

Pay Back Period = 1 + 4/5.40 = 1.74 years

Profitability Index = PV of Cash Inflows / PV of Cash Outflows = 6.34 / 4 = 1.59

Investment Option 4:

Year Cash Flows in millions PV factor @ 6% Total PV in millions Cumulative Cash Flow
0 (5.00) 1.0 (5.00) (5.00)
1 5.20 0.9434 4.91 0.20
2 1.60 0.8900 1.42 1.80
NPV $ 1.33

Pay Back Period = 1 + 0.20/5.20 = 1.04 years

Profitability Index = PV of Cash Inflows / PV of Cash Outflows = 6.33 / 5 = 1.27

Based on the Net Present Value, Ginny should choose Investment option 2 as its profitability index is the highest and has low payback period

The Solution to Question 1 part iii)

Ginny should eliminate investment option 3 as its payback period is high for similar profitability index return options.